OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83? -200 -197 -194 -191

OpenStudy (anonymous):

The approach is similar: find the common difference by subtracting the two terms and dividing; so d = (a17 - a1)/16 = (-83-13)/16 = -6 Now a35 = a1 + 34d = 13 + 34(-6) = 13 - 204 = -191